OpenStudy (anonymous):

Find the slope and y-intercept of each line. 4x+14y= 21

OpenStudy (dumbcow):

solve for y

OpenStudy (dumbcow):

once you get it in y= mx +b form m -> slope b-> y-intercept

Directrix (directrix):

4x+14y= 21 ========== 14y = 21 - 4x y = 21/14 - (4/14) x y = 3/2 -(2/7) x y = -(2/7) x + 3/2 slope: -2/7; y-intercept: 3/2 http://www.wolframalpha.com/input/?i=4x%2B14y%3D+21